3. Exactly how were the MPC and MPS in Table 21-4 computed? Illustrate by calculating MPC and MPS between
points A and B. Explain why it must always be true that MPC+MPS = 1
1
Expert's answer
2021-10-11T09:57:41-0400

Solution:

The MPC has been calculated as follows:

MPC = "\frac{Change\; in\; Consumption\; expenditure}{Change\; in \;Disposable\; Income \;(after\; taxes}"



The MPS has been calculated as follows:

MPS ="\frac{Change\; in\; Savings}{Change\; in \;Disposable\; Income \;(after\; taxes}"


This is illustrated as follows for points A and B:

Change in Consumption expenditure between points A and B = (25,000 – 24,200) = 800

Change in Disposable Income (after taxes) between points A and B = (25,000 – 24,000) = 1000

MPC = "\frac{800}{1000} = 0.80"


Change in Net Savings between points A and B = (0 – (-200)) = 200

Change in Disposable Income (after taxes) between points A and B = (25,000 – 24,000) = 1000

MPS = "\frac{200}{1000} = 0.20"

 

It is always true that MPC + MPS must equal to one since MPC is the fraction of the change in income spent, therefore, the fraction not spent must be saved and this is the MPS. Since the denominator is the total change in income, the sum of MPC and MPS is equal to one.